OpenStudy (anonymous):

x^@-2x=-2 answwers will be in comments

OpenStudy (anonymous):

a) x=-1+-i sqrt{2} b) x=1+- i sqrt{2} c) x=-1+-i d)x=1+-i

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

its, \(x^2-2x=-2\) right ?

OpenStudy (anonymous):

yes that person that was trying to help me on the other question confused me you helped lol but yes that is the equation

hartnn (hartnn):

write it as \(x^2-2x+2=0\) Compare ^ quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] \[ \\ \sqrt{b^2-4ac}=...?\] then the two roots of x are: \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

a=1 b=-2 c=2

hartnn (hartnn):

correct , b^2-4ac =... ?

OpenStudy (anonymous):

-2^2-4(1)(2)=-4

hartnn (hartnn):

good! :) so, now just plug in the values in the formula

OpenStudy (anonymous):

..... watttttt

hartnn (hartnn):

\(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

oh

hartnn (hartnn):

can you solve firther ?

hartnn (hartnn):

\(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}=\frac{2 \pm \sqrt{-4}}{2}=\frac{2 \pm 2\sqrt{-1}}{2}\) you know \(\sqrt{-1}=i\) ?

OpenStudy (anonymous):

no not really this is where i get confused and yes i know the sqrt of-1 is i

hartnn (hartnn):

\(\huge \frac{2 \pm 2\sqrt{-1}}{2}=\frac{2 \pm 2i}{2}=\dfrac{2}{2}\pm\dfrac{2i}{2}\) got this ?

OpenStudy (anonymous):

yes so it should be d right

hartnn (hartnn):

yes :D

OpenStudy (anonymous):

yay im still gonna need your help on a few more questions

hartnn (hartnn):

sorry, i have to leave, i need to go somewhere urgently....

OpenStudy (anonymous):

oh ok